I need to know the diagonal distance of a 50 ft square please
Answer by nerdybill(7384)

I need to know the diagonal distance of a 50 ft square please
.
Applying Pythagorean theorem:
Let d = diagonal
then
d^2 = 50^2 + 50^2
d^2 = 2500 + 2500
d^2 = 5000
d = 70.71 feet